OpenStudy (anonymous):

Which expression is equivalent to 27 + 45? 8(3 + 6) 8(19 + 37) 9(3 + 5) 9(18 + 36)

OpenStudy (anonymous):

this is distributive property. So if you multiply the outside times the inside, it will be 9*3 = 27 and 9*5= 45 Its 9(3+5)

OpenStudy (anonymous):

27+45= 72 so it would be 8(3+6) because 3+6=9 and 8*9=72

OpenStudy (anonymous):

8(3 + 6) this is the one that i think

OpenStudy (anonymous):

@sethRB26DETT but think about it, it's not asking for the answer. It's asking for an equation that is equivalent

OpenStudy (anonymous):

IT'S A TRAP!!!!!!!!!

OpenStudy (anonymous):

both are equivalent @ste.cano

OpenStudy (rhystic):

no ste.cano is correct. I misread. it wants an expression that equates to 27+45 not the answer. so 9(3+5) is correct because 9(3+5) = 27+45

OpenStudy (anonymous):

am i right

OpenStudy (rhystic):

8(3+6) = 24+ 48 which is incorrect.

OpenStudy (anonymous):

exactly but its asking for you to get the same answer as the equation so its basically asking you to simplify

OpenStudy (anonymous):

??

OpenStudy (anonymous):

will medal

OpenStudy (anonymous):

@Rhystic please explain how you got 8(3+6)=24+48

OpenStudy (anonymous):

they both give you the same equation but the question is asking you to simplify in order to get the same equation And they multiplied 8*3 and 8*6 which is correct

OpenStudy (anonymous):

answer*

OpenStudy (anonymous):

??? so what you just said is right right

OpenStudy (rhystic):

matching the expression would be more correct which 9(3+5)=27+45=72 becaues 8(3+6) = 24+ 48 which is not the same as 27+45 although both equal 72.

OpenStudy (anonymous):

All you have to do in problems like this is multiply the outside number * the inside numbers inside the parenthesis

OpenStudy (anonymous):

okay so you distribute the 8 into the parentheses i thought it was solving the expression

OpenStudy (rhystic):

9(3+5) = 9*3 + 9*5 = 27+45 ; 8(3+6) = 8*3+8*6 = 24 + 48

OpenStudy (anonymous):

@sethRB26DETT All it was asking for was to get the same equation not the same answer